AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 110     Entry time: Wed Oct 6 09:47:30 2021
Author: CB RSS 
Subject: Neg activation tests -2 
Attempted activating 2xUHV1400 connected to top C4, channel #3 with P 300 W (450 C?), t rise 30 min, t hold 90
min as yesterday

Power reach 300 W, hold. Temperature increased by 5C. OK.

Activated top C4, channel #4 using same set of parameters. Internal temperature increases significantly, up to
140C and still increasing.

Decreased power on channel #2 to 100 W (down from 200 W yesterday). Should correspond to 350 C or so.
Temperature still increasing.
Decreased power on channels #3 and #4 to 250 W (425 C?). Temperature stops increasing. Around 155C on kapton
bundle. OK for kapton, detectors would have broken.
Baking code detected temperature on internals >150 C and stopped baking for safety. Temperatures do not change -
heating provided by NEGs more than enough to maintain current status.

Decreased power on channel #2 to 50W (250 C?). Temperature on internals decreases slightly. No major effect. 
This is likely because the UHV1400 corresponding to #2 are far from the detectors, while #3 and #4 are closer.

Decided to complete activation of #3 and #4 using 250 W. Set both to 50C after completion. Temperature of
internals decreases quickly.

Re-activated baking code when temperature on all internals <150 C. Setpoint 85C on RE72-1 and 95 on -2.

Now waiting for internals to cool down. Will activate remaining 3 sets (2+2+3 UHV1400) later today with P 200 W
(400 C?) t rise 30 min, t hold 90 min.
We know the set of 3 UHV1400 is definitely close to the kapton bundle. P 300 W (100x3) for that set.
ELOG V3.1.4-unknown